# Plasmon Enhanced Spectroscopy and Photocatalysis

### v1

@@ 4 Conclusions
In this study, we have presented two interrelated projects that test the ability of TD-DFTB to describe the influence of plasmon excitation on two physical phenomena: surface-enhanced Raman scattering (SERS) and plasmon-enhanced charge transfer. In the first, we calculated SERS spectra for Au-pyridine complexes, where we found enhancements due to both chemical and plasmon excitation effects that are consistent with earlier TD-DFT studies.TDDFT-Schatz-2006 (19) The enhancement factors for chemical effect are around 3-15 depending on the site of the adsorption. For the electromagnetic effect, the enhancement is on the order of for the V-complex (for adsorption on one of the tips) whereas for the S-complex (for adsorption on one of the triangle faces). In the second project, we studied plasmon-driven charge transfer across particle-molecule junctions after laser excitation of the particle near the plasmon energy at 2.5 eV. We found substantial electron transfer from particle to molecule for a Au-C complex leading to the formation of a negative molecular ion, C on the particle. This observation is consistent with a previous experimental result that shows evidence of electron transfer from particle to molecule and the formation of negative molecular ions for a fullerene derivative.SERS-PCBM-Schatz-2019 (33) Noteworthy in this result is that we did not see other photoprocesses in these simulations of laser pulses with intensities that are close to threshold, including C-C bond breakage, fragmentation of the particle, or desorption of the fullerene from the particle. The ability of the DFTB method to simulate SERS and charge transfer problems opens several new possibilities in the plasmonics field.

### v3

@@ 4 Conclusions
In this study, we have presented two interrelated projects that test the ability of TD-DFTB to describe the influence of plasmon excitation on two physical phenomena: surface-enhanced Raman scattering (SERS) and plasmon-enhanced charge transfer. In the first, we calculated SERS spectra for Au120-pyridine complexes, where we found enhancements due to both chemical and plasmon excitation effects that are consistent with earlier TD-DFT studies.TDDFT-Schatz-2006 The enhancement factors for chemical effect are around 3-15 depending on the site of the adsorption. For the electromagnetic effect, the enhancement is on the order of for the V-complex (for adsorption on one of the tips) whereas for the S-complex (for adsorption on one of the triangle faces). In the second project, we studied plasmon-driven charge transfer across particle-molecule junctions after laser excitation of the particle near the plasmon energy at 2.5 eV. We found substantial electron transfer from particle to molecule for a Au120-C60 complex leading to the formation of a negative molecular ion, C on the particle. This observation is consistent with a previous experimental result that shows evidence of electron transfer from particle to molecule and the formation of negative molecular ions for a fullerene derivative.SERS-PCBM-Schatz-2019 Noteworthy in this result is that we did not see other photoprocesses in these simulations of laser pulses with intensities that are close to threshold, including C-C bond breakage, fragmentation of the particle, or desorption of the fullerene from the particle. The ability of the DFTB method to simulate SERS and charge transfer problems opens several new possibilities in the plasmonics field.